PART 1 A ball of mass 2.5 kg hangs at the end of a string 1.1...

PART 1

A ball of mass 2.5 kg hangs at the end of a string 1.1 m in length from the ceiling of a room. The height of the room is 3.6 m. What is the gravitational potential energy of the ball relative to the ceiling?

Tries 0/2

What is the gravitational potential energy of the ball relative to the floor?

Tries 0/2

What is the gravitational potential energy (in J) of the ball relative to a point at the same elevation as the ball?

PART 2

Forces of 70 N at 80 degrees, and 20 N at an angle of 230 degrees, measured counter-clockwise from the positive x-axis, act on an object.

What is the magnitude of the resultant force (in Newtons)?

Tries 0/2

What is the angle of the resultant force with respect to x-axis? (positive values only)
Answer #1

Solution)

Part 1)

Given, mass m=2.5 kg

We know, P.E=mgh

PE=2.5*9.8*(-1.1)=- 26.95 J

======

Relative to the floor

PE= mgh'

PE=2.5*9.8*(3.6-1.1) = 61.25 J

======

Same elevation as ball

PE=mgh =2.5*9.8*0= 0J

=======